OpenStudy (highschoolmom2010):

use eulers formula to find number of vertices in each polyhedron. 5 faces, 1 rectangle, 4 triangles

ganeshie8 (ganeshie8):

Euler's formula :- V - E + F = 2

ganeshie8 (ganeshie8):

faces, F = 5

OpenStudy (highschoolmom2010):

my book says it is F+V=E+2

OpenStudy (anonymous):

Either way, they both would equal the same answer

OpenStudy (highschoolmom2010):

i dont exactly understand what to do:

OpenStudy (highschoolmom2010):

5 faces 8 edges 5+V=8+2 5+V=10 -5 -5 V=5

ganeshie8 (ganeshie8):

you nailed it ! good job !!

OpenStudy (highschoolmom2010):

alright thanks @ganeshie8 @Ryan2416

ganeshie8 (ganeshie8):

np :)